OpenStudy (anonymous):

csc^2(x) + sec^2(x) = csc^2(x)sec^2(x)

OpenStudy (anonymous):

what is your question?

OpenStudy (anonymous):

prove the trig identity so I have to make the LS = RS

OpenStudy (raden):

sometimes, we can starting from the right side. we knowed that csc = 1/sin and sec = 1/cos so, csc^2(x) sec^2(x) = 1/sin^2(x) * 1/cos^2(x) = 1/(sin^2x cos^2 x) remmber this identity : 1 = sin^2x + cos^2x therefore, 1/(sin^2x cos^2 x) = (sin^2(x) + cos^2(x))/(sin^2x cos^2 x) splite into 2 parts, get = sin^2x/(sin^2x cos^2 x) + cos^2x/(sin^2x cos^2 x) = 1/cos^2x + 1/sin^2(x) = sec^2(x) + csc^2(x)

OpenStudy (anonymous):

would you always want to split one side if its like A + B / AB? so it would then be A / AB + B / AB ? did i write that right?

OpenStudy (anonymous):

LHS\[\frac{ 1 }{ \sin^2x }+\frac{ 1 }{ \cos^2x }=\frac{ 1 }{ sin^2x* cos^2x }\] \[\frac{ \sin^2+\cos^2x }{ \sin^2x* \cos^2x}=\frac{ 1 }{ \sin^2x *\cos^2x}\] \[LHS=\frac{ 1 }{ \sin^2x *\cos^2x }=\frac{ 1 }{ \sin^2x *\cos^2x }=RHS\]

OpenStudy (raden):

yes, then cancels wich similar terms

OpenStudy (raden):

* which

OpenStudy (anonymous):

alright thanks again!

OpenStudy (raden):

welcome